# Detecting Near-Duplicate Face Images

### v1

@@ Abstract
Near-duplicate images are often generated when applying repeated photometric and geometric transformations that produce imperceptible variants of the original image. Consequently, a deluge of near-duplicates can be circulated online posing copyright infringement concerns. The concerns are more severe when biometric data is altered through such nuanced transformations. In this work, we address the challenge of near-duplicate detection in face images by, firstly, identifying the original image from a set of near-duplicates and, secondly, deducing the relationship between the original image and the near-duplicates. We construct a tree-like structure, called an Image Phylogeny Tree (IPT) using a graph-theoretic approach to estimate the relationship, i.e., determine the sequence in which they have been generated. We further extend our method to create an ensemble of IPTs known as Image Phylogeny Forests (IPFs). We rigorously evaluate our method to demonstrate robustness across other modalities, unseen transformations by latest generative models and IPT configurations, thereby significantly advancing the state-of-the-art performance by $\sim 42\%$ on IPF reconstruction accuracy.
@@ 1 Introduction
Figure 1: Examples of near-duplicates available online illustrating celebrity face images modified through (top row) Photoshop, (middle row) FaceApp and (bottom row) MakeApp filters. Image courtesy: https://nypost.com/2019/07/16/faceapp-challenge-celebrities-reveal-how-theyll-look-as-seniors/
@@
Image editing tools can seamlessly generate near-duplicates by sequentially applying transformations to an original image. Photos, especially face images, appearing on social networking sites can be subjected to image enhancement techniques for enhancing visual appeal [1, 2]. Celebrity magazines sometimes use ‘airbrushing’ to glamorize advertisements and this has drawn the attention of lawmakers. [^1] https://www.telegraph.co.uk/news/2017/09/30/photoshopped-images-come-warning-new-french-law/ Photo-editing filters and generative models such as generative adversarial networks (GANs) and diffusion models can be used to alter the appearance of any individual by modifying attributes such as age and hairstyle [3, 4]. See Fig. 1. Multiple digital modifications can result in an explosive number of near-duplicates that end up online with minimal regulation. In the wake of such unmitigated data duplication, provenance analysis has piqued the interest of academic and government entities [5]. It aids in tracing the evolution of the edited images and plays an integral role in media forensics. An image phylogeny tree (IPT) provides a visual assessment of the evolution and relationship between near-duplicates (see Figs. 2 and 3) and is used for their detection [6, 7]. IPT reconstruction begins with a set of near-duplicate images as input and generates a hierarchical structure as output. It consists of a root node (the original image) and directed edges (the relationship between the near-duplicates); the edges are directed from the parent node (source image) to the child node (transformed image). Although IPT consists of a single root node, there can be a situation involving multiple root nodes. An example can be a face image of an individual in the same scene but with different poses, expressions or captured using different cameras. In this case, transforming each image repeatedly, but independently of the others, will result in multiple IPTs. Each original image will then serve as a separate root, and each of them will span a distinct IPT. A collection of such IPTs will constitute an Image Phylogeny Forest (IPF) [8]. See Fig. 4.
@@
Our Method: Existing methods [6, 9, 10, 11, 12, 13, 14] focus on images depicting natural scenes and generic objects but face images pose a difficult challenge due to their inherent diversity and heterogeneity in terms of age, gender, ethnicity, expression, illumination, pose, and other attributes. Additionally, face images possess biometric utility that should not be hampered during processing. Therefore, we formulate the phylogeny tree construction as a graph-based framework where each image acts as a node, and the directed edges indicate the relationship between the nodes. A graph-based approach is well-suited for image phylogeny as it allows both pairwise and global analyses among the nodes. The nodes (images) are represented via their respective features, while the edges (interactions) are represented through an adjacency matrix that captures the relationship between them. We begin with an initial adjacency matrix and iteratively refine it until we arrive at the desired configuration resulting in the IPT. We employ a convolutional graph neural network that utilizes the feature values of the nodes to learn the interactions that best capture the relationship between the nodes. Further, we consider multiple IPTs and devise a simple technique to construct the image phylogeny forest (IPF) with high degrees of accuracy. The contributions of this work are as follows.
@@
1. We develop a technique that combines a graph neural network with sensor noise pattern (SPN) features to examine local and global interaction between the nodes, i.e., a set of near-duplicates. The graph network serves as a node embedding module that identifies the hierarchical position of each near-duplicate node in an IPT. The SPN features perform link prediction between each pair of nodes embedded by the graph network. The node embedding and link prediction modules work in tandem to perform complementary local and global analyses to generate the IPT. We extend our method to create an IPF using a locally-scaled spectral clustering method that identifies disjoint IPTs.
@@ 3 Proposed Method
Next, we proceed to the second phase of our IPF pipeline, i.e., recovering the phylogeny structure within each cluster. Our method couples graph theory with sensor noise pattern features to explore macroscopic (global) and microscopic (pairwise) analysis of the images. Consider the near-duplicates as nodes or vertices arranged in a high-dimensional space such that they are related to each other because they are derived from each other in a specific order. The order in which they have been modified can then be represented as a ‘walk’ from one node to the other. IPT reconstruction can therefore be considered as determining the correct sequence of walks between the nodes which will indicate the global relationship between a set of near-duplicates. We use node embedding to derive this global relationship. Node embedding converts a high-dimensional input (an image) to a low-dimensional output (a scalar value). The scalar value generated by the node embedding should indicate the global relationship of a node. But the challenge is how can a single scalar value represent a global property? We formulated the problem as follows: we will use node embedding to deduce the depth label (a scalar value) of a node (image). The depth label of an image (say, $k$) in an IPT signifies the number of ancestors ($k-1$) for that image, with respect to the root node located at depth=1. For example, an image at depth=2 has only 1 ancestor, i.e., the root node, while an image at depth=3 has 2 ancestors. The depth labels of the images in an IPT provide how the images are globally related with each other. However, the depth labels cannot determine the relationship between the nodes at successive depths. For example, there can be multiple nodes with depth labels $s$ but only one of them will be the immediate parent of node associated with depth label $t$ (where, $t=s+1$). Therefore, we use a link prediction module to identify the existence of links between nodes located at depth $k$ and depths $>k$ by performing a pairwise microscopic analysis. See details of the node embedding module in Sec. 3.2 and the link prediction module in Sec. 3.3.
@@ 3.1 Locally-scaled spectral clustering
1. Construct a symmetric similarity matrix $\bm{S}$ and its corresponding adjacency matrix $\bm{A}$ from the data points $\bm{X}$.
@@
We apply the local-scaling on the feature space ($f(\bm{X})$) instead of directly applying it on the data space ($\bm{X}$) to allow aggregation of multiple features for accurate clustering. We selected three features in this work: (i) pixel intensity, (ii) sensor noise pattern features, particularly Enhanced Photo Response Non-Uniformity (PRNU) [39], and (iii) face descriptors. Pixel intensity and PRNU features account for the variation in the acquisition settings and devices. Face descriptors [^2] We can always substitute face descriptors with generic image descriptors for images depicting natural scenes. account for variation in pose and expression. We use VGGFace descriptors [40] in this work. The resultant similarity matrix now becomes,
@@
Here, $\bm{D}_{F}(i,j)=1-\frac{\bm{F}_{i}\bm{F}_{j}}{\bm{F}_{i}^{2}\bm{F}_{j}^{2}}$, indicates the cosine distance between the face descriptors $F(\cdot)$; $\bm{D}_{N}(i,j)=1-\frac{\bm{N}_{i}\bm{N}_{j}}{\bm{N}_{i}^{2}\bm{N}_{j}^{2}}$, indicates the cosine distance between the PRNU $N(\cdot)$ features; $\bm{D}_{P}(i,j)=1-\frac{\bm{P}_{i}\bm{P}_{j}}{\bm{P}_{i}^{2}\bm{P}_{j}^{2}}$, indicates the cosine distance between the pixel intensity features $P(\cdot)$. The terms in the denominator are locally-scaled bandwidths computed using univariate kernel density estimate of the respective features. $\hat{p}_{F}(\cdot)$ corresponds to kernel density estimate (KDE) computed from face descriptors, $\hat{p}_{F}(\cdot)=\frac{1}{nh}\sum_{i=1}^{n}K\bigg{(}\frac{F(\cdot)-F(x_{i})}{h}\bigg{)}$, where $n$ is the number of data points, $h$ is the bandwidth equal to the interval between all the points in $x$ and $K$ is a normal kernel. Similarly, $\hat{p}_{N}(\cdot)$ denotes KDE computed from the PRNU features, and $\hat{p}_{P}(\cdot)$ denotes KDE computed from the pixel intensity features. Next, we compute binarized similarity matrix $\bm{S}_{B}$ using median of each row as threshold, such that $\bm{S}_{B}(i,j)=1$ if $\bm{S}(i,j)>median(\bm{S}(i,:))$, and $\bm{S}_{B}(i,j)=0$, otherwise. Then we follow the steps pertaining to conventional spectral clustering, such as computation of the degree matrix, Laplacian matrix, and eigen decomposition. The degree matrix is computed as $\bm{Deg}(i,i)=\sum_{j=1}^{Q}\bm{S}_{B}(i,j)$. The normalized Laplacian is computed as $\bm{L}=\bm{Deg}^{-\frac{1}{2}}\bm{S}_{B}\bm{Deg}^{-\frac{1}{2}}$. Eigen value decomposition of the Laplacian matrix results in eigen vectors ($\bm{U}$) and eigen values ($\Lambda$). We apply a threshold $\eta$ (selected using a validation set) to obtain a subset of eigen values, say $k$ smallest eigen values such that, $\{\Lambda_{k}\subset\Lambda|\Lambda_{k}<\eta\}$. The corresponding eigen vectors become $\bm{U}_{k}=[\bm{u}_{1},\bm{u}_{2},\cdots,\bm{u}_{k}]$. Finally, k-means is applied to the rows of $\bm{U}_{k}$ to get cluster assignments, $\mathscr{C}\in\mathbb{R}^{Q\times 1}$ for each image, such that $\mathscr{C}=$ k-means$(\bm{U}_{k},k)$, where, $k$ is the number of clusters. See Algo. 1.
@@
8      7 Compute distance values between each pair of images ($\bm{I}_{i},\bm{I}_{j}$) for face descriptors, PRNU features and pixel features, respectively (we used cosine distance)                                                        $\bm{D}_{F}(i,j)=1-\frac{\bm{F}_{i}\bm{F}_{j}}{\bm{F}_{i}^{2}\bm{F}_{j}^{2}}$                                                        $\bm{D}_{N}(i,j)=1-\frac{\bm{N}_{i}\bm{N}_{j}}{\bm{N}_{i}^{2}\bm{N}_{j}^{2}}$                                                        $\bm{D}_{P}(i,j)=1-\frac{\bm{P}_{i}\bm{P}_{j}}{\bm{P}_{i}^{2}\bm{P}_{j}^{2}}$                                                        Compute symmetric similarity matrix $\bm{S}$:                                                        $\bm{S}(i,j)=\exp-\bigg{(}\frac{\bm{D}_{F}(i,j)^{2}}{\hat{p}_{F}(i)\hat{p}_{F}(j)}+\frac{\bm{D}_{N}(i,j)^{2}}{\hat{p}_{N}(i)\hat{p}_{N}(j)}+\frac{\bm{D}_{P}(i,j)^{2}}{\hat{p}_{P}(i)\hat{p}_{P}(j)}\bigg{)}$
9until $i,j\leq Q$;
10Compute binarized similarity matrix $\bm{S}_{B}$ using median of each row as threshold
11Compute diagonal degree matrix $\bm{Deg}$
12Compute normalized Laplacian $\bm{L}$
13Perform eigen value decomposition to obtain eigen vectors ($\bm{U}$) and eigen values ($\Lambda$) from Laplacian
14Select thereshold $\eta$ to select $k$ smallest eigen vectors $\Lambda_{k}$: $\bm{U}_{k}=[\bm{u}_{1},\bm{u}_{2},\cdots,\bm{u}_{k}]$
15Apply k-means clustering on the rows of $\bm{U}_{k}$ to get cluster assignments $\mathscr{C}$ for each image
@@
Figure 5: Outline of the two-step method used in this work. In the first step, the node embedding module, typically a graph neural network, $f\bm{(X,A)}$ accepts $\bm{X}$: pixel intensity values of each image in the IPT and $\bm{A}$: an initial adjacency matrix as a pair of inputs. The node embedding module produces a vector of depth labels corresponding to each IPT configuration. In the second step, the link prediction module accepts $g\bm{(l,N)}$, $\bm{l}$: depth labels from the first module, and $\bm{N}$: sensor noise pattern (PRNU) features computed from the near-duplicates as inputs. The link prediction module produces the reconstructed image phylogeny tree (IPT).

### v2

@@ Abstract
Near-duplicate images are often generated when applying repeated photometric and geometric transformations that produce imperceptible variants of the original image. Consequently, a deluge of near-duplicates can be circulated online posing copyright infringement concerns. The concerns are more severe when biometric data is altered through such nuanced transformations. In this work, we address the challenge of near-duplicate detection in face images by, firstly, identifying the original image from a set of near-duplicates and, secondly, deducing the relationship between the original image and the near-duplicates. We construct a tree-like structure, called an Image Phylogeny Tree (IPT) using a graph-theoretic approach to estimate the relationship, i.e., determine the sequence in which they have been generated. We further extend our method to create an ensemble of IPTs known as Image Phylogeny Forests (IPFs). We rigorously evaluate our method to demonstrate robustness across other modalities, unseen transformations by latest generative models and IPT configurations, thereby significantly advancing the state-of-the-art performance by $\sim 42\%$ on IPF reconstruction accuracy. Our code is publicly available at https://github.com/sudban3089/DetectingNear-Duplicates.
@@ 1 Introduction
Figure 1: Examples of near-duplicates available online illustrating celebrity face images modified through (top row) Photoshop, (middle row) FaceApp and (bottom row) MakeApp filters. Image courtesy [1].
Image editing tools can seamlessly generate near-duplicates by sequentially applying transformations to an original image. Photos, especially face images, appearing on social networking sites can be subjected to image enhancement techniques for enhancing visual appeal [2, 3]. Celebrity magazines sometimes use ‘airbrushing’ to glamorize advertisements and this has drawn the attention of lawmakers. [^1] https://www.telegraph.co.uk/news/2017/09/30/photoshopped-images-come-warning-new-french-law/ Photo-editing filters and generative models such as generative adversarial networks (GANs) and diffusion models can be used to alter the appearance of any individual by modifying facial and demographic attributes such as age and hairstyle [4, 5]. See Fig. 1.
@@
Problem Definition. Multiple digital modifications can result in an explosive number of near-duplicates that end up online with minimal regulation. In the wake of such unmitigated data duplication, provenance analysis has piqued the interest of academic and government entities [6]. It aids in tracing the evolution of the edited images and plays an integral role in media forensics. An image phylogeny tree (IPT) provides a visual assessment of the evolution and relationship between near-duplicates (see Figs. 2 and 3) and is used for their detection [7, 8]. IPT reconstruction begins with a set of near-duplicate images as input and generates a hierarchical structure as output. It consists of a root node (the original image) and directed edges (the relationship between the near-duplicates); the edges are directed from the parent node (source image) to the child node (transformed image). Although IPT consists of a single root node, there can be a situation involving multiple root nodes. An example can be a face image of an individual in the same scene but with different poses, expressions or captured using different cameras. In this case, transforming each image repeatedly, but independently of the others, will result in multiple IPTs. Each original image will then serve as a separate root, and each of them will span a distinct IPT. A collection of such IPTs will constitute an Image Phylogeny Forest (IPF) [9].
@@
Our Method: Existing methods [7, 10, 11, 12, 13, 14, 15] focus on images depicting natural scenes and generic objects but face images pose a difficult challenge due to their inherent diversity and heterogeneity in terms of age, gender, ethnicity, expression, illumination, pose, and other attributes. Additionally, face images possess biometric utility that should not be hampered during processing. Therefore, we formulate the phylogeny tree construction as a graph-based framework where each image acts as a node, and the directed edges indicate the relationship between the nodes. A graph-based approach is well-suited for image phylogeny as it allows both pairwise and global analyses among the nodes. The nodes (images) are represented via their respective features, while the edges (interactions) are represented through an adjacency matrix that captures the relationship between them. We begin with an initial adjacency matrix and iteratively refine it until we arrive at the desired configuration resulting in the IPT. We employ a convolutional graph neural network that utilizes the feature values of the nodes to learn the interactions that best capture the relationship between the nodes. Further, we consider multiple IPTs and devise a simple technique to construct the image phylogeny forest (IPF) with high degrees of accuracy. See Fig. 4. The contributions of this work are as follows.
@@
1. We develop a technique that combines a graph neural network with sensor pattern noise (SPN) features to examine local and global interaction between the nodes, i.e., a set of near-duplicates. The graph network serves as a node embedding module that identifies the hierarchical position of each near-duplicate node in an IPT. The SPN features perform link prediction between each pair of nodes embedded by the graph network. The node embedding and link prediction modules work in tandem to perform complementary local and global analyses to generate the IPT. We extend our method to create an IPF using a locally-scaled spectral clustering method that identifies disjoint IPTs.
@@ 3 Proposed Method
Next, we proceed to the second phase of our IPF pipeline, i.e., recovering the phylogeny structure within each cluster. Our method couples graph theory with sensor pattern noise features to explore macroscopic (global) and microscopic (pairwise) analysis of the images. Consider the near-duplicates as nodes or vertices arranged in a high-dimensional space such that they are related to each other because they are derived from each other in a specific order. The order in which they have been modified can then be represented as a ‘walk’ from one node to the other. IPT reconstruction can therefore be considered as determining the correct sequence of walks between the nodes which will indicate the global relationship between a set of near-duplicates. We use node embedding to derive this global relationship. Node embedding converts a high-dimensional input (an image) to a low-dimensional output (a scalar value). The scalar value generated by the node embedding should indicate the global relationship of a node. But the challenge is how can a single scalar value represent a global property? We formulated the problem as follows: we will use node embedding to deduce the depth label (a scalar value) of a node (image). The depth label of an image (say, $k$) in an IPT signifies the number of ancestors ($k-1$) for that image, with respect to the root node located at depth=1. For example, an image at depth=2 has only 1 ancestor, i.e., the root node, while an image at depth=3 has 2 ancestors. The depth labels of the images in an IPT provide how the images are globally related with each other. However, the depth labels cannot determine the relationship between the nodes at successive depths. For example, there can be multiple nodes with depth labels $s$ but only one of them will be the immediate parent of node associated with depth label $t$ (where, $t=s+1$). Therefore, we use a link prediction module to identify the existence of links between nodes located at depth $k$ and depths $>k$ by performing a pairwise microscopic analysis. See details of the node embedding module in Sec. 3.2 and the link prediction module in Sec. 3.3.
@@ 3.1 Locally-scaled spectral clustering
1. Construct a symmetric similarity matrix $\bm{S}$ and its corresponding adjacency matrix $\bm{A}$ from $Q$ data points $\bm{X}$.
@@
We apply the local-scaling on the feature space ($f(\bm{X})$) instead of directly applying it on the data space ($\bm{X}$) to allow aggregation of multiple features for accurate clustering. We selected three features in this work: (i) pixel intensity, (ii) sensor pattern noise features, particularly Enhanced Photo Response Non-Uniformity (PRNU) [40], and (iii) face descriptors. Pixel intensity and PRNU features account for the variation in the acquisition settings and devices. Face descriptors [^2] We can always substitute face descriptors with generic image descriptors for images depicting natural scenes. account for variation in pose and expression. We use VGGFace descriptors [41] in this work. The resultant similarity matrix now becomes,
@@
The cosine similarity between face descriptors (vectors) computed from a pair of face images $(i,j)$, i.e., $\textbf{F}_{i}$ and $\textbf{F}_{j}$ is $cosinesim(\textbf{F}_{i},\textbf{F}_{j})=\frac{\textbf{F}_{i}^{T}\textbf{F}_{j}}{\lVert\textbf{F}_{i}\rVert\lVert\textbf{F}_{j}\rVert}.$ Considering the cosine similarity to vary between $[-1,1]$, we first normalize it to $[0,1]$ via $0.5\times(cosinesim+1)$. Then we use the normalized value to compute a distance measure as $\textbf{D}_{F}(i,j)=1-\frac{\textbf{F}_{i}^{T}\textbf{F}_{j}}{\lVert\textbf{F}_{i}\rVert\lVert\textbf{F}_{j}\rVert}.$ Similarly, cosine distance between pixel features, $P(\cdot)$, can be expressed as $\textbf{D}_{P}(i,j)=1-\frac{\textbf{P}_{i}^{T}\textbf{P}_{j}}{\lVert\textbf{P}_{i}\rVert\lVert\textbf{P}_{j}\rVert}.$ Finally, cosine distance between PRNU-based sensor noise features, $N(\cdot)$, can be expressed as $\textbf{D}_{N}(i,j)=1-\frac{\textbf{N}_{i}^{T}\textbf{N}_{j}}{\lVert\textbf{N}_{i}\rVert\lVert\textbf{N}_{j}\rVert}.$
The terms in the denominator are locally-scaled bandwidths computed using univariate kernel density estimate of the respective features. $\hat{p}_{F}(\cdot)$ corresponds to kernel density estimate (KDE) computed from face descriptors, $\hat{p}_{F}(\cdot)=\frac{1}{nb}\sum_{i=1}^{n}h\bigg{(}\frac{F(\cdot)-F(x_{i})}{b}\bigg{)}$, where $n$ is the number of data points, $b$ is the bandwidth equal to the interval between all the points in $x$ and $h$ is a normal kernel. Similarly, $\hat{p}_{N}(\cdot)$ denotes KDE computed from the PRNU features, and $\hat{p}_{P}(\cdot)$ denotes KDE computed from the pixel intensity features. Next, we compute binarized similarity matrix $\bm{S}_{B}$ using median of each row as threshold, such that $\bm{S}_{B}(i,j)=1$ if $\bm{S}(i,j)>median(\bm{S}(i,:))$, and $\bm{S}_{B}(i,j)=0$, otherwise. Then we follow the steps pertaining to conventional spectral clustering, such as computation of the degree matrix, Laplacian matrix, and eigen decomposition. The diagonal degree matrix for $Q$ nodes is computed as $\bm{Deg}(i,i)=\sum_{j=1}^{Q}\bm{S}_{B}(i,j)$. The normalized Laplacian is computed as $\bm{L}=\bm{Deg}^{-\frac{1}{2}}\bm{S}_{B}\bm{Deg}^{-\frac{1}{2}}$. Eigen value decomposition of the Laplacian matrix results in eigen vectors ($\bm{U}$) and eigen values ($\Lambda$). We apply a threshold $\eta$ (selected using a validation set) to obtain a subset of eigen values, say $k$ smallest eigen values such that, $\{\Lambda_{k}\subset\Lambda|\Lambda_{k}<\eta\}$. The corresponding eigen vectors become $\bm{U}_{k}=[\bm{u}_{1},\bm{u}_{2},\cdots,\bm{u}_{k}]$. Finally, k-means is applied to the rows of $\bm{U}_{k}$ to get cluster assignments, $\mathscr{C}\in\mathbb{R}^{Q\times 1}$ for each image, such that $\mathscr{C}=$ k-means$(\bm{U}_{k},k)$, where, $k$ is the number of clusters. See Algo. 1.
@@
7       Compute distance values between each pair of images ($\bm{I}_{i},\bm{I}_{j}$) for face descriptors, PRNU features and pixel features, respectively (we used cosine distance)                                                        $\textbf{D}_{F}(i,j)=1-\frac{\textbf{F}_{i}^{T}\textbf{F}_{j}}{\lVert\textbf{F}_{i}\rVert\lVert\textbf{F}_{j}\rVert}$
8       $\textbf{D}_{N}(i,j)=1-\frac{\textbf{N}_{i}^{T}\textbf{N}_{j}}{\lVert\textbf{N}_{i}\rVert\lVert\textbf{N}_{j}\rVert}$
10      9 $\textbf{D}_{P}(i,j)=1-\frac{\textbf{P}_{i}^{T}\textbf{P}_{j}}{\lVert\textbf{P}_{i}\rVert\lVert\textbf{P}_{j}\rVert}$                                                        Compute symmetric similarity matrix $\bm{S}$:                                                        $\bm{S}(i,j)=\exp-\bigg{(}\frac{\bm{D}_{F}(i,j)^{2}}{\hat{p}_{F}(i)\hat{p}_{F}(j)}+\frac{\bm{D}_{N}(i,j)^{2}}{\hat{p}_{N}(i)\hat{p}_{N}(j)}+\frac{\bm{D}_{P}(i,j)^{2}}{\hat{p}_{P}(i)\hat{p}_{P}(j)}\bigg{)}$
11until $i,j\leq Q$;
12Compute binarized similarity matrix $\bm{S}_{B}$ using median of each row as threshold
13Compute diagonal degree matrix $\bm{Deg}$
14Compute normalized Laplacian $\bm{L}$
15Perform eigen value decomposition to obtain eigen vectors ($\bm{U}$) and eigen values ($\Lambda$) from Laplacian
16Select thereshold $\eta$ to select $k$ smallest eigen vectors $\Lambda_{k}$: $\bm{U}_{k}=[\bm{u}_{1},\bm{u}_{2},\cdots,\bm{u}_{k}]$
17Apply k-means clustering on the rows of $\bm{U}_{k}$ to get cluster assignments $\mathscr{C}$ for each image
@@
Figure 5: Outline of the two-step method used in this work. In the first step, the node embedding module, i.e., a graph neural network, $f\bm{(X,A)}$ accepts $\bm{X}$: features from each image in the IPT; features can be pixel intensity or convolutional features, and $\bm{A}$: an initial adjacency matrix as a pair of inputs. During training, the initial adjacency matrix denotes the ground truth relationship between the images in the IPT, while during testing, the initial adjacency matrix is populated by comparing distance between images features against a threshold to determine presence of edge $(=1)$ or absence of edge $(=0)$. The node embedding module produces a vector of depth labels corresponding to each IPT configuration. In the second step, the link prediction module accepts $g\bm{(l,N)}$, $\bm{l}$: depth labels from the first module, and $\bm{N}$: sensor pattern noise (PRNU) features computed from the near-duplicates as inputs. The link prediction module produces the reconstructed image phylogeny tree (IPT).